What is the linear system of 2x+y=2,2x+3y=13 using substitution?
Answer by pwac(253)

2x+y=2------(1)
2x+3y=13----(2)
Rearrange(1)to get y=2-2x.Substitute this into (2)
2x+3(2-2x)=13 Multiply bracket by 3
2x+6-6x=13 Rearrange to get:
-4x=7 Divide both sides by -4
x=-1.75
Now substitute this value of x in equation (1)
2(-1.75)+y=2
-3.5+y=2
y=5.5
Check by using values of x and y in (1)
2(-1.75)+5.5=2 correct
Pete